### v2

@@ IV Asymptotics of $k^{*}(n)$
We now return to the phenomenon in the two-dimensional Fermi-Hubbard model (Fig. 1) in which the value $\hat{R}(k)$ saturates for some $k^{*}<n$. In this case we see numerically that $k^{*}(n)$ scales linearly with $n$ by repeating the experiment on $n=40,80,$ and $120$ qubits. Thus, while $k$-commutativity leads to a practical reduction in measurement complexity, the asymptotics are still the same. In this section we consider the question of which Hamiltonians, if any, exhibit a $k^{*}(n)$ which scales sublinearly with $n$. In particular, we consider several physically inspired models and find examples for which $k^{*}(n)=O(1)$ and $k^{*}(n)=O(\sqrt{n})$, the latter case being the same Fermi-Hubbard Hamiltonian in Fig. 1 but using the Jordan-Wigner encoding instead of the Bravyi-Kitaev encoding.
Before analyzing the Fermi-Hubbard model, we first consider if there are physically-inspired classes of $n$-qubit Hamiltonions for which $k^{*}(n)=O(1)$. This indeed is true by considering the one-dimensional transverse field Ising model
@@
Here, the $Z$ type terms in $H_{n}$ $k$-commute for any $k$, and similarly for the $X$ type terms, thus there are exactly two groups for any $k$ and $k^{*}(n)=O(1)$. Other cases for which $k^{*}(n)=O(1)$ include qubit-based simulation of low-energy Bose-Hubbard models [elstner1999dynamics, somma2003quantum, sawaya2022mat2qubit] and quantum molecular vibrational Hamiltonians [mcardle2019digital, sawaya2020connectivity, ollitrault2020hardware] for which there is a bounded degree of connectivity between vibrational modes. This is because (unlike the case of fermionic commutation) bosonic commutation leads to qubit Hamiltonians with bounded locality.
We now consider the spinless Fermi-Hubbard Hamiltonian (5) under the Jordan-Wigner encoding, starting for simplicity in one dimension. In the Jordan-Wigner encoding, we map the fermion operators $a_{i},a_{i}^{\dagger}$ to qubit operators via
$a_{i}$ $\rightarrow Z_{1}\otimes\cdots\otimes Z_{i-1}\otimes S^{-}_{i}$ (8)
$a_{i}^{\dagger}$ $\rightarrow Z_{1}\otimes\cdots\otimes Z_{i-1}\otimes S^{+}_{i}$ (9)
where $S^{+}=|1\rangle\langle 0|$ and $S^{-}=(S^{+})^{\dagger}$. The hopping terms take the form
$\begin{split}a_{i+1}^{\dagger}a_{i}&=(Z_{1}\otimes\cdots\otimes Z_{i-1}\otimes Z_{i}\otimes S^{+}_{i+1})\\ &\times(Z_{1}\otimes\cdots\otimes Z_{i-1}\otimes S^{-}_{i}\otimes I_{i+1})\\ &=Z_{i}S^{-}_{i}\otimes S^{+}_{i+1}\\ &=S^{-}_{i}\otimes S^{+}_{i+1}\\ \end{split}$ (10)
and similarly
$\begin{split}a_{i-1}^{\dagger}a_{i}&=(Z_{1}\otimes\cdots\otimes Z_{i-2}\otimes S^{+}_{i-1}\otimes I_{i})\\ &\times(Z_{1}\otimes\cdots\otimes Z_{i-2}\otimes Z_{i-1}\otimes S^{-}_{i})\\ &=S^{+}_{i-1}Z_{i-1}\otimes S^{-}_{i}\\ &=S^{+}_{i-1}\otimes S^{-}_{i}\\ \end{split}$ (11)
We then express the raising and lower operators of the spins in terms of Pauli operators
$S^{\pm}=\frac{1}{2}(X\pm iY)$ (12)
giving
$a_{i+1}^{\dagger}a_{i}=\frac{1}{2}(X_{i}-iY_{i})\otimes\frac{1}{2}(X_{i+1}+iY_{i+1})$ (13)
and
$a_{i-1}^{\dagger}a_{i}=\frac{1}{2}(X_{i-1}+iY_{i-1})\otimes\frac{1}{2}(X_{i}-iY_{i}).$ (14)
The strings of the form $\{X_{0}X_{1},X_{1}X_{2},\ldots\}$ are qubit-wise commuting, as are $\{Y_{0}Y_{1},Y_{1}Y_{2},\ldots\}$. Observe that
$[X\otimes Y\otimes I,I\otimes Y\otimes X]=0.$ (15)
Thus we can build a $k$-commuting set $\{X_{0}Y_{1},Y_{1}X_{2},X_{2}Y_{3},\ldots\}$ (for any $k$). Another one can be built by switching $X$ and $Y$. The hopping terms in the Hamiltonian can be sorted into four commuting groups with just $k=1$, regardless of lattice depth.
Next consider Coulomb terms of the form $a_{i}^{\dagger}a_{j}^{\dagger}a_{i}a_{j}$. In one dimension, the nearest neighbor points are $j=i+1$ and $j=i-1$. Much like with the hopping terms above, these Coulomb terms will only act on neighboring qubits.
$\begin{split}a_{i}^{\dagger}a_{i+1}^{\dagger}a_{i}a_{i+1}&=(S^{+}_{i}Z_{i}S^{-}_{i}Z_{i})(S^{+}_{i+1}S^{-}_{i+1})\\ &=-S^{+}_{i}S^{-}_{i}\otimes S^{+}_{i+1}S^{-}_{i+1}\\ &=-|1\rangle_{i}\prescript{}{i}{\langle 1|}\otimes|1\rangle_{i+1}\prescript{}{i+1}{\langle 1|}\\ &=-\frac{1}{2}(I_{i}-Z_{i})\otimes\frac{1}{2}(I_{i+1}-Z_{i+1})\\ \end{split}$ (16)
This forms a set of qubit-wise commuting Pauli strings.
Finally, consider an on-site potential term $a_{i}^{\dagger}a_{i}$. Again the strings of $Z$ operators cancel, leaving only operators acting on qubit $i$. These can be sorted into $k$-commuting groups for any $k$. Thus, the whole Fermi-Hubbard Hamiltonian can be can be sorted into groups with only $k=1$, regardless of the size of the lattice. However, when $k=2$, we can sort the terms that go like $X_{i}X_{i+1}$ and $Y_{i}Y_{i+1}$ into a single group, unlike in the $k=1$ case where these terms must go into different groups. Splitting a group of operators into two groups means that $\hat{R}$ must be less after the splitting [Crawford_Straaten_Wang_Parks_Campbell_Brierley_2021], so having these $XX$ and $YY$ terms in the same group results in a higher $\hat{R}$. This can be seen in Fig. 3 on the edges where $M=1$ or $N=1$ and $k^{*}(n)=2$. Note that this demonstrates an interesting example where $k^{*}(n)=O(1)$ but $k^{*}(n)\neq 1$ as in the previously-mentioned Hamiltonians such as the Ising model.
Now consider the two-dimensional case, with lattice size $M\times N$. We will use row-major order, i.e. the first $M$ qubits correspond to the first row, the next $M$ to the second row, etc. Hopping terms between two rows now take the form
$a^{\dagger}_{i+M}a_{i}=S_{i}^{-}\otimes Z_{i+1}\otimes\ldots\otimes Z_{i+M-1}\otimes S_{i+M}^{+}$ (17)
and
$a^{\dagger}_{i}a_{i+M}=S_{i}^{+}\otimes Z_{i+1}\otimes\ldots\otimes Z_{i+M-1}\otimes S_{i+M}^{-}$ (18)
Expanding $S^{\pm}$ in terms of Pauli operators, we get Pauli strings of the forms
$X\otimes Z\otimes\ldots\otimes Z\otimes X,$
$X\otimes Z\otimes\ldots\otimes Z\otimes Y,$
$Y\otimes Z\otimes\ldots\otimes Z\otimes X,$
and
$Y\otimes Z\otimes\ldots\otimes Z\otimes Y$
acting on contiguous blocks of $M+1$ qubits. Due to the fact that $[X,Z]\neq 0$ and $[Y,Z]\neq 0$, two Pauli strings must act on completely separate blocks of qubits, and $k\geq M$, in order for two strings to commute. For example, the strings $XZZY$ and $IXZZY$ do not commute, no matter the block size, as do $XZZY$ and $IIXZZY$. The cutoff value $k^{*}$ after which $\hat{R}$ does not increase should thus be $M+1$. Figure 3 shows the values of $k^{*}(M,N)$ computed numerically for various lattices sizes and matches the result $k^{*}(M,N)=M+1$, or in the case of square lattices that $k^{*}(n)=O(\sqrt{n})$.
Figure 3: The values of $k^{*}(n)=k^{*}(M,N)$ for the spinless, aperiodic Fermi-Hubbard Hamiltonian (5) under the Jordan-Wigner transformation on an $M\times N$ lattice. These values, computed numerically, match the derivation that $k^{*}(M,N)=M+1$ for $M,N>1$, or for square grids that $k^{*}(n)=O(\sqrt{n})$.
The case of the Fermi-Hubbard model in one and two dimensions shows why it is useful to generalize commutation checks beyond the two extreme cases of qubit-wise commutativity and (full) commutativity previously considered in the literature. The energy of the one-dimensional model can be measured using only single-qubit rotations, since all operators are qubit-wise commuting. Previous methods work well in this case. For the two-dimensional model on an $M\times N$ lattice, we identify a cutoff value $k^{*}(M,N)=M+1$ after which $\hat{R}$ does not increase. The number of shots needed to measure the energy in this case is thus the same if one uses (fully) $NM$-commuting or $(M+1)$-commuting strings. Since fewer gates are needed to diagonalize the $(M+1)$-commuting sets, the notion of $k$-commutativity (here with $k=M+1$) saves on the total resources (shots and circuit depth) needed to measure the energy.
Figure 4: The values of $k^{*}(n)$ for random Hamiltonians. For each $n$, a random Hamiltonian $H_{n}$ with $n$ terms is constructed by sampling, for each term, a number of Paulis $t$ from the exponential distribution $p(t)=\frac{1}{w}\exp(-wt)$. Three different values for $w$ are chosen: constant $w=2$, logarithmic $w=\ln n$, and linear $w=n/2$. After sampling a number of Paulis $t$, each Pauli ($X$, $Y$, or $Z$) and each index $i=0,...,n-1$ is chosen uniformly at random to form the term. This process repeats for all terms. The value of $k^{*}$ is determined as in Fig. 1. Error bars show the standard deviation over fifty random $H_{n}$. As can be seen in the plot of $k^{*}(n)$, the scaling is linear for $w=O(n)$ and constant for $w=O(1)$.
@@
As a final example, we numerically compute the value of $k^{*}(n)$ for random $n$-qubit Hamiltonians in one dimension. To generate these Hamiltonians $H_{n}$, we set the number of terms equal to $n$. For each term, a number $t$ of Paulis is sampled from the exponential distribution $p(t)=\frac{1}{w}\exp(-wt)$ with parameter $w$. Then, $t$ Paulis $X$, $Y$, $Z$ are chosen uniformly at random, and the qubits that they act on are chosen uniformly at random. Once the Hamiltonian $H_{n}$ is constructed, $k^{*}(n)$ is numerically computed, and this process repeats for fifty random instances. The results are shown in Fig. 4 for three values of $w$, namely constant $w=2$, logarithmic $w=\ln n$, and linear $w=n/2$. As can be seen in Fig. 4, we observe that the $k^{*}(n)$ is linear for $w=O(n)$ and constant for $w=O(1)$. We thus conjecture that random $n$-qubit Hamiltonians with a linear number of terms of constant weight are another family of Hamiltonians for which $k^{*}(n)=O(1)$, though we leave a rigorous analysis of this conjecture for future work. One potential application of this conjecture is in error mitigation via quantum subspace expansion [McClean_Jiang_Rubin_Babbush_Neven_2020]. In this framework, one measures the expectation value of an observable projected into the codespace of an error correcting code. By using error correcting codes for which $k^{*}(n)=O(1)$, for example random codes with constant weight stabilizers, one could potentially measure error-mitigated observables with measurement complexity constant or sublinear in $n$.
@@ V Conclusion
In this work we have introduced a simple notion of commutativity for Pauli strings that considers blocks of size $k\in\mathbb{N}$, and motivated this notion by the practical application of measurement reduction when evaluating expectation values in quantum circuits. In addition to defining $k$-commutativity we have proved several properties of the relation including the size of $k$-commuting sets and the depth of diagonalization circuits. Numerically, we have shown a reduction in the measurement complexity of the two-dimensional spinless Fermi-Hubbard model and various molecular Hamiltonians. Finally, we built on an observation in the Fermi-Hubbard model that measurement complexity saturates for some $k^{*}<n$ and characterized the asymptotic measurement complexity of $k$-commutativity for several families of $n$-qubit Hamiltonians.
@@
Code and data availability Software to reproduce numerical results can be found in this Colab.
Acknowledgements This work was supported by Wellcome Leap as part of the Quantum for Bio Program. RS would like to thank the Institute for Pure and Applied Mathematics for being generous hosts and for providing a great environment during the period when this work was completed.
@@ Appendix A Lower bounds on the Clifford gate complexity for diagonalizing commuting Paulis
Since each element of the Clifford group $\mathcal{C}_{n}$ is a commutation preserving group isomorphism of $\mathcal{P}_{n}$, we conclude that the set :=$\mathcal{B}:=\{P_{1},P_{2},\dots,P_{n},Q_{1},Q_{2},\dots,Q_{n}\}$ is also independent, and together with $iI$ generates $\mathcal{P}_{n}$.
@@
since every gate can be chosen from the gate set in $n(n-1)+2n+1$ ways. First let us fix any such Clifford unitary $U\in\mathcal{U}$. We know by Lemma 1 that modulo phase factors, all the Paulis diagonalized upon conjugation by $U$ belongs to a commuting subgroup of size $2^{n}$. Let us call this subgroup $\mathcal{S}$. As a symplectic vector space, $\mathcal{S}$ is a $\mathbb{F}_{2}$-vector space of dimension $n$. Thus, the number of ways of choosing $r\leq n$ independent elements of $\mathcal{S}$, is the same as choosing $r$ linearly independent elements of $\mathbb{F}_{2}^{n}$, and by Lemma 3 this count is given by
@@
Figure 5: The number of groups and $\hat{R}$ metrics obtained for sorted insertion (left panel) and random insertion (right panel) on the $n=10$ qubit $C_{2}$ molecule [hamlib]. For random insertion, error bars show the standard deviation over $100$ trials. Note that the number of groups is comparable for both algorithms but the values for $\hat{R}$ are substantially different. In particular, $\hat{R}$ is significantly lower for the random insertion algorithm. This, as well as the relatively large variance in $\hat{R}$ for the random insertion algorithm, provides another example showing that $\hat{R}$ is a more appropriate metric than the number of groups.
Appendix C Additional numerical results
Here we include additional numerical results, in particular the analogue Fig. 2 using the Jordan-Wigner encoding is shown in Fig. 6, and the analogue of Fig. 3 using the Bravyi-Kitaev encoding is shown in Fig. 7.
For the molecular Hamiltonians (Fig. 6), we see comparable results to Fig. 2. The behavior that $\hat{R}$ increases as $k$ increases is still seen and the vertical axis is on the same scale, but for some molecules $\hat{R}$ seems to increase more quickly under than Bravyi-Kitaev encoding than the Jordan-Wigner encoding, notably $O_{2}$ and $B_{2}$.
For the Fermi-Hubbard model (Fig. 7), the behavior of $k^{*}(n)$ under the Bravyi-Kitaev encoding is qualitatively similar to the case of the Jordan-Wigner encoding (Fig. 3), however the pattern is less regular and the values are generally higher.
Figure 6: The same as Fig. 2 but using the Jordan-Wigner encoding instead of the Bravyi-Kitaev encoding.
Figure 7: The same as Fig. 3 but using the Bravyi-Kitaev encoding instead of the Jordan-Wigner encoding.

### v3

@@ IV Asymptotics of $k^{*}(n)$
We now return to the phenomenon of the Bacon-Shor Hamiltonian as shown in Fig. 1. Here, we see the value of $\hat{R}(k)$ reaches a maximum for $k^{*}=O(\sqrt{n})$. In this section we prove this scaling and consider the question of which other Hamiltonians, if any, exhibit a $k^{*}(n)$ which scales sublinearly with $n$.
We first prove that $k^{*}(n)=O(\sqrt{n})$ for the Bacon-Shor Hamiltonian (5) on a square lattice ($M=N$) of $n=M\times N$ qubits. The intuition for why this is true is because stabilizer elements for the Bacon-Shor code row-wise and column-wise commute, formally proved in the following lemma.
Lemma 1:
Any pair of elements $S$ and $S^{\prime}$ from the stabilizer group for the Bacon-Shor code on $M\times N$ qubits $M$-commute column-wise.
Proof.
Let $P\in\mathcal{P}_{N}$ be a Pauli string and let $R\subseteq\{1,2,\dots,N\}$ be a subset of qubit indices. The restriction of $P$ to subset $R$ is given by $P|_{R}\equiv\bigotimes_{i\in R}P_{i}$. The restriction of a group $\mathcal{G}$ to a subset $R$ is the group generated by the generators of $\mathcal{G}$ restricted to the same subset. For the Pauli group, this restriction is always a group. We define the support of $P$ to be the set of qubits where $P$ acts non-trivially, $\text{supp}(P):=\left\{i|P_{i}\neq I\right\}$. For a group $\mathcal{G}$, $\text{supp}\left(\mathcal{G}\right)$ is the union of the supports of the elements of $\mathcal{G}$.
Consider two Bacon-Shor stabilizer groups $\mathcal{S}$ and $\mathcal{T}$ for systems of size $M\times N$ and $M\times\left(N+1\right)$ respectively. $\mathcal{S}$ and $\mathcal{T}$ are both Abelian and their group elements commute with any choice of Bacon-Shor logical such as $\bar{Z}_{N-1}=\bigotimes_{k=0}^{M-1}Z_{k,N-1}$. Observe $\mathcal{T}|_{\text{supp}(S)}=\mathcal{S}\cup\left(\bigotimes_{k=0}^{M-1}\,Z_{k,N-1}Z_{k,N}\right)\Big|_{\text{supp}(\mathcal{S})}=\mathcal{S}\cup\bar{Z}_{N-1}$. For any two stabilizers $t_{1},t_{2}\in\mathcal{T}$, we have $0=[t_{1},t_{2}]=[t_{1}|_{\text{supp}(\mathcal{S})},t_{2}|_{\text{supp}(\mathcal{S})}]+[t_{1}|_{\text{supp}(\mathcal{S})^{c}},t_{2}|_{\text{supp}(\mathcal{S})^{c}}]=0+[t_{1}|_{\text{column N}},t_{2}|_{\text{column N}}]$. Since the weight of the stabilizer group restricted to a column is $M$, the stabilizers $M$-commute column-wise. ∎
Note that the same argument applies when exchanging rows and columns, thus we have the following corollary.
Corollary 2:
Any pair of elements $S$ and $S^{\prime}$ from the stabilizer group for the Bacon-Shor code on $M\times N$ qubits $N$-commute row-wise.
Now, consider a square $M=N$ lattice of $n$ qubits. Since all coefficients in the Bacon-Shor Hamiltonian are one, $\hat{R}$ is maximized by minimizing the number of groups. The above lemma shows that all terms in the Hamiltonian can be placed into one group by taking $k*=M=N=\sqrt{n}$, thus we have proved the following.
Theorem 5:
For the Bacon-Shor Hamiltonian (5) on a square lattice of $n$ qubits, $k^{*}(n)=O(\sqrt{n})$.
Figure 4: The values of $k^{*}(n)$ — i.e., the value of $1\leq k\leq n$ in $k$-commuting for which measurement reduction saturates — for $n$-qubit random Hamiltonians. For each $n$, a random Hamiltonian $H_{n}$ with $n$ terms is constructed by sampling, for each term, a number of Paulis $t\in\{0,...,n\}$ from the exponential probability density function $p(t)=\frac{1}{w}\exp(-wt)$ (rounded to the nearest integer). Three different values for $w$ are chosen: constant $w=2$, logarithmic $w=\ln n$, and linear $w=n/2$. After sampling a number of Paulis $t$, each Pauli ($X$, $Y$, or $Z$) and each index i=0,,n1i=0,...,n-1 is chosen uniformly at random to form the term. This process repeats for all terms. The value of $k^{*}$ is determined as in Fig. 1. Error bars show the standard deviation over fifty random $H_{n}$. As can be seen in the plot of $k^{*}(n)$, the scaling is linear for $w=O(n)$ and constant for $w=O(1)$.
Because the Bacon-Shor Hamiltonian exhibits $k^{*}(n)=O(\sqrt{n})$ scaling, it is natural to ask if there are other Hamiltonian families which exhibit similar or even better scaling. After brief consideration, it is easy to see that there are physically-inspired classes of $n$-qubit Hamiltonians for which $k^{*}(n)=O(1)$. This indeed is true by considering the one-dimensional transverse field Ising model
@@
Here, the $Z$ type terms in $H_{n}$ $k$-commute for any $k$, and similarly for the $X$ type terms, thus there are exactly two groups for any $k$ and $k^{*}(n)=O(1)$. (It’s easy to see that there cannot be fewer than two groups, and because the coefficients for all $Z$-type terms are identical and all coefficients for all $X$-type terms are identical, it follows that $\hat{R}$ is maximized by this grouping.)
Another less obvious family of Hamiltonians for which $k^{*}(n)=O(1)$ is the Bose-Hubbard model [somma2003quantum, sawaya2022mat2qubit, bahrami2024particle]
$H=\sum_{ij}t_{ij}(b_{i}^{\dagger}b_{j}+h.c.)+g\sum_{i}n_{i}(n_{i}-1)$ (8)
@@
where $b_{i}^{\dagger}$ ($b_{i}$) is the bosonic creation (annihilation) operator for mode/site $i$, $t_{ij}$ is a hermitian matrix giving the kinetic energy of a boson moving from site $i$ to site $j$, $g$ is the site energy, and $n_{i}:=b_{i}^{\dagger}b_{i}$ is the number operator. The intuition for why $k^{*}(n)=O(1)$ for this model is that, unlike the case of fermionic commutation, bosonic commutation leads to qubit Hamiltonians with bounded locality. This same observation holds for quantum molecular vibrational Hamiltonians [mcardle2019digital, sawaya2020connectivity, ollitrault2020hardware] for which there is a bounded degree of connectivity between vibrational modes. Here, we consider only the case of the Bose-Hubbard Hamiltonian, and discuss this in more detail.
We consider hardcore bosons in which there can be at most one boson per lattice site. (This can be caused by strong repulsive interactions.) Since there is at most one boson per site, we can represent states in the hardcore boson model by using the qubit states $|0\rangle$ and $|1\rangle$ to represent Fock states with zero and one bosons, respectively. With this representation, we can represent $b_{j}^{\dagger}$ by the qubit operator $|1\rangle\langle 0|=\frac{1}{2}(X_{j}-iY_{j})$, and $b_{j}$ by $|0\rangle\langle 1|=\frac{1}{2}(X_{j}+iY_{j})$. Thus, the hardcore boson qubit Hamiltonian is, in one dimension,
$H=\frac{t}{2}\sum_{i}(X_{i}X_{i+1}+Y_{i}Y_{i+1})+2g\sum_{j}(I-Z_{j}).$
In the $k=1$ case, all of the $Z_{j}$ terms can be put into the same group. Similarly, all of the $X_{i}X_{i+1}$ terms $1$-commute with each other, as do the $Y_{i}Y_{i+1}$ terms. Thus, similar to the TFIM, this model exhibits $k^{*}=O(1)$. In Appendix C, we include numerical calculations of $k^{*}$ without assuming hardcore bosons. These results suggest that $k^{*}(n)=O(1)$ for the Bose-Hubbard model in general, which we here leave as a conjecture.
@@
As a final example, we numerically compute the value of $k^{*}(n)$ for random $n$-qubit Hamiltonians in one dimension. To generate these Hamiltonians $H_{n}$, we set the number of terms equal to $n$. For each term, a number $t$ of Paulis is sampled from the exponential distribution $p(t)=\frac{1}{w}\exp(-wt)$ with parameter $w$. Then, $t$ Paulis $X$, $Y$, $Z$ are chosen uniformly at random, and the qubits that they act on are chosen uniformly at random. Once the Hamiltonian $H_{n}$ is constructed, $k^{*}(n)$ is numerically computed, and this process repeats for fifty random instances. The results are shown in Fig. 4 for three values of $w$, namely constant $w=2$, logarithmic $w=\ln n$, and linear $w=n/2$. As can be seen in Fig. 4, we observe that the $k^{*}(n)$ is markedly linear for $w=O(n)$ and $w=O(\log n)$. Up to $n=100$, the case for $w=O(1)$ is less clear. However, based on the numerics and intuition from the previously-studied case of the Bacon-Shor Hamiltonian, we conjecture that random $n$-qubit Hamiltonians with a linear number of terms of constant weight are another family of Hamiltonians for which $k^{*}(n)=O(1)$, though we leave a rigorous analysis of this conjecture for future work. One potential application of this conjecture is in error mitigation via quantum subspace expansion [McClean_Jiang_Rubin_Babbush_Neven_2020]. In this framework, one measures the expectation value of an observable projected into the codespace of an error correcting code. By using error correcting codes for which $k^{*}(n)=O(1)$, for example random codes with constant weight stabilizers, one could potentially measure error-mitigated observables with measurement complexity constant or sublinear in $n$.
@@ V Conclusion
In this work we have introduced a simple notion of commutativity for Pauli strings that considers blocks of size $k\in\mathbb{N}$, and motivated this notion by the practical application of measurement reduction when evaluating expectation values in quantum circuits. In addition to defining $k$-commutativity we have proved several properties of the relation including the size of $k$-commuting sets and the depth of diagonalization circuits. Numerically, we have shown a reduction in the measurement complexity of various Hamiltonians including those from chemical, condensed matter, and error correction models. Finally, we built on an observation in the Bacon-Shor Hamiltonian that measurement complexity saturates for some $k^{*}<n$ and characterized the asymptotic measurement complexity of $k$-commutativity for several families of $n$-qubit Hamiltonians.
@@
Code and data availability Software to reproduce results can be found at github.com/rmlarose/kcommute.
Acknowledgments This work was supported by the Wellcome Leap Quantum for Bio Program. RS thanks the Institute for Pure and Applied Mathematics for being generous hosts and for providing a great environment during the period when this work was completed.
@@ Appendix A Lower bounds on the Clifford gate complexity for diagonalizing commuting Paulis
Since each element of the Clifford group $\mathcal{C}_{n}$ is a commutation preserving group isomorphism of $\mathcal{P}_{n}$, we conclude that the set $\mathcal{B}:=\{P_{1},P_{2},\dots,P_{n},Q_{1},Q_{2},\dots,Q_{n}\}$ is also independent, and together with $iI$ generates $\mathcal{P}_{n}$.
@@
since every gate can be chosen from the gate set in n(n1)+2n+1n(n-1)+2n+1 ways. First let us fix any such Clifford unitary $U\in\mathcal{U}$. We know by Lemma 2 that modulo phase factors, all the Paulis diagonalized upon conjugation by $U$ belongs to a commuting subgroup of size $2^{n}$. Let us call this subgroup $\mathcal{S}$. As a symplectic vector space, $\mathcal{S}$ is a $\mathbb{F}_{2}$-vector space of dimension $n$. Thus, the number of ways of choosing $r\leq n$ independent elements of $\mathcal{S}$, is the same as choosing $r$ linearly independent elements of $\mathbb{F}_{2}^{n}$, and by Lemma 4 this count is given by
@@
Figure 5: The number of groups and $\hat{R}$ metrics obtained for sorted insertion (left panel) and random insertion (right panel) on the $n=10$ qubit $C_{2}$ molecule [hamlib]. For random insertion, error bars show the standard deviation over 100100 trials. Note that the number of groups is comparable for both algorithms but the values for $\hat{R}$ are substantially different. In particular, $\hat{R}$ is significantly lower for the random insertion algorithm. This, as well as the relatively large variance in $\hat{R}$ for the random insertion algorithm, provides another example showing that $\hat{R}$ is a more appropriate metric than the number of groups when considering measurement reduction.
Appendix C Numerics for the Bose-Hubbard model
Here, we consider the Bose-Hubbard Hamiltonian (8) for the case where there can be more than one boson per site. The Bose-Hubbard models in HamLib [hamlib] have examples of both unary and binary mappings between Fock states of the bosons and computational basis states of the qubits. W numerically compute $k^{*}$ for several one-dimensional Bose-Hubbard models from HamLib. Table 2 shows results for the unary encoding, and Table 3 shows results for the binary encoding. Our numerical experiments show that the unary encoding gives a value of $k^{*}=4$ for all lattice sizes we investigated, while in general for the binary encoding we see that $k^{*}=n$. Similarly, the unary case shows a greater ratio of the maximal $\hat{R}$ value to the minimal $\hat{R}$ value. The unary encoding is thus more efficient in terms of the number of shots needed for measuring the Hamiltonian. Note, however, that the number of qubits needed to encode the Hamiltonian is much larger, $O(r)$ for the unary encoding, and $\log(r)$ for the binary encoding, where $r$ is the maximal number of bosons at each site. A detailed view of the algorithm’s behavior is given by the left sub-plot in Fig. 6, which shows $\hat{R}$ and the number of groups as a function of $k$ for a unary-encoded Bose-Hubbard model. A similar plot is shown for a two-dimensional case of the same Hamiltonian in the right sub-plot of Fig. 6. Numerics for this case show that the maximal value of $\hat{R}$ is achieved when $k=4$. However, the the ratio of the maximal and minimal $\hat{R}$ values is only 1.05. The performance of our algorithm for the one-dimensional case seems much better than in two dimensions.
Figure 6: Left: $\hat{R}$ (orange) and number of groups (blue) vs. $k$ for a unary-encoded, one-dimensional Bose-Hubbard model of lattice size $l=12$, with at most 4 bosons per site. Right: $\hat{R}$ (orange) and number of groups (blue) vs. $k$ for a unary-encoded, two-dimensional Bose-Hubbard model of lattice size $5\times 5$, with at most 4 bosons per site.
$l$ $n$ $k^{*}$ $\hat{R}_{\text{max}}/\hat{R}_{\text{min}}$
4 16 4 1.65
10 40 4 1.33
12 48 4 1.66
16 64 4 1.15
22 88 4 1.07
Table 2: Lattice dimension $l$, number of qubits $n$, the value $k^{*}$ that maximizes $\hat{R}$, and the ratio of the maximal to minimal $\hat{R}$ values for the one-dimensional Bose-Hubbard model in the unary encoding.
$l$ $n$ $k^{*}$ $\hat{R}_{\text{max}}/\hat{R}_{\text{min}}$
4 8 2 1.21
10 20 20 1.33
12 24 24 1.32
16 32 32 1.14
22 44 44 1.05
Table 3: Lattice dimension $l$, number of qubits $n$, the value $k^{*}$ that maximizes $\hat{R}$, and the ratio of the maximal to minimal values of $\hat{R}$ for the one-dimensional Bose-Hubbard model in the binary encoding.